OpenStudy (anonymous):

can someone help me solve this one. Best Rentals charges a daily fee plus a mileage fee for renting its cars. Barney was charged $123.00 for 3 days and 300 miles, while Mary was charged $216.00 for 5 days and 600 miles. What does Best Rental charge per day, for mileage?

OpenStudy (anonymous):

put \[x=\text{ price per day}\] \[y=\text{ mileage fee}\] then you get two equations \[3x+300y=123\] and \[5x+600y=216\] and now solve for x and y

OpenStudy (anonymous):

BarneyFee = $123 = 3*DayFee + 300*miles*milageFee MaryFee = $216 = 5*DayFee + 600*miles*milageFee

OpenStudy (anonymous):

i like touchymath way. but you still have to solve

OpenStudy (anonymous):

i would multiply \[3x+300y=123\] by -2 to get \[-6x-600y=-246\] then add to \[5x+600y=216\] to get \[-x=-30\] or \[x=30\]so $30 per day

OpenStudy (anonymous):

The steps to the solution are in the attached file DayFee = $30/day MileFee = $0.11/mile

OpenStudy (anonymous):

Thanks touchyMath and Satellite73

OpenStudy (anonymous):

@touchymath what program is that?

OpenStudy (anonymous):

math101...and its been a while since i had math subject

OpenStudy (anonymous):

good luck!

OpenStudy (anonymous):

thanks